# RainPro-8: An Efficient Deep Learning Model to Estimate Rainfall Probabilities Over 8 Hours

### v1

@@ Abstract
We present a deep learning model for high-resolution probabilistic precipitation forecasting over an 8-hour horizon in Europe, overcoming the limitations of radar-only deep learning models with short forecast lead times. Our model efficiently integrates multiple data sources - including radar, satellite, and physics-based numerical weather prediction (NWP) - while capturing long-range interactions, resulting in accurate forecasts with robust uncertainty quantification through consistent probabilistic maps. Featuring a compact architecture, it enables more efficient training and faster inference than existing models. Extensive experiments demonstrate that our model surpasses current operational NWP systems, extrapolation-based methods, and deep-learning nowcasting models, setting a new standard for high-resolution precipitation forecasting in Europe, ensuring a balance between accuracy, interpretability, and computational efficiency.
@@ 1 Introduction
Recent advances in artificial intelligence, especially in computer vision and natural language processing, have generated significant interest in deep learning for weather forecasting (Rasp et al., 2024; An et al., 2024). Deep learning excels at handling complex, large-scale, high-dimensional datasets, making it well-suited for capturing intricate, nonlinear patterns in spatio-temporal systems (Manzhu Yu and Li, 2024). Although deep learning has achieved remarkable success in both nowcasting (Gao et al., 2024b; Gong et al., 2024) and medium-range forecasting (Lam et al., 2023; Price et al., 2024), significant challenges remain. Deep learning models for nowcasting are often limited to very short lead times (up to two hours). In contrast, medium-range models, which predict broader atmospheric dynamics for up to 10 days, typically operate at coarser resolutions and are influenced by precipitation-specific biases in the training datasets (Lavers et al., 2022). As a result, they struggle to capture small-scale precipitation features, like local showers, often leading to the exclusion of precipitation forecasts in medium-range models (Lam et al., 2023).
@@
This work addresses the challenge of forecasting precipitation for up to 8 hours at high spatio-temporal resolutions of 2 km per pixel and 10-minute intervals, bridging the gap between nowcasting and medium-range forecasting. Forecasting over an 8-hour horizon is critical for timely predictions that help mitigate risks like flooding and optimize resource management in sectors such as agriculture, energy, and transportation. This task is particularly difficult due to the stochastic and sparse nature of precipitation, especially over extended lead times. Designing a deep learning model for this task requires addressing skewed and intermittent precipitation distributions, incorporating multi-sensor data, and ensuring robust uncertainty quantification.
@@
Most existing deep learning-based nowcasting methods focus on generating radar-like precipitation forecasts (Ravuri et al., 2021; Gao et al., 2024b), but these become increasingly challenging and less accurate with longer lead times. Instead, probabilistic models emphasize uncertainty estimation and provide a more accurate and practical representation of rainfall at varying intensities. The MetNet family of models (Sønderby et al., 2020; Espeholt et al., 2022; Andrychowicz et al., 2023) represents state-of-the-art deep learning-based systems for probabilistic precipitation forecasting. These models produce high-resolution forecasts for 8-24 hours in the United States, outperforming operational NWP systems. They achieve efficient probabilistic forecasting using cross-entropy loss on precipitation bins, requiring only one forward pass rather than an ensemble. However, this approach neglects the ordinality between bins. In addition, they rely on lead-time conditioning, where the model generates forecasts for only one lead time at a time. While this approach is fully parallelizable across multiple GPUs, it significantly increases inference computational demands. Moreover, training MetNet requires substantial computational resources and is limited to the United States.
@@
We propose RainPro-8, an efficient deep learning model to estimate rainfall probabilities over 8 hours in Europe by determining the probability of different levels of precipitation at a given location and time. RainPro-8 is an efficient model based on MetNet-3 (Andrychowicz et al., 2023) that uses less than 20% of MetNet-3’s training parameters to achieve 8-hour high-resolution precipitation forecasting across Europe — a region characterized by diverse climates, complex terrain, and highly variable precipitation dynamics (Ehmele et al., 2020). We also redesign the training objective to explicitly account for the ordinality between precipitation bins. Lastly, we demonstrate that it is possible to generate all lead times simultaneously by downweighing later lead times during training, greatly enhancing inference efficiency and improving temporal consistency. Our model demonstrates substantial improvements over traditional precipitation forecasting methods, including NWP models and extrapolation-based techniques, as well as deep-learning nowcasting models, across various rain intensities for the next 8 hours.
Our main contributions are as follows:
@@
  • We introduce an efficient model architecture and training strategy that learns from multi-source data with varying temporal and spatial resolutions, accounts for the ordinality of precipitation classes, and outputs all lead times simultaneously.
  • We evaluate our approach on a European dataset we created specifically for 8-hour precipitation nowcasting and demonstrate that our method outperforms existing operational methods by 65%, as well as deep-learning nowcasting models.
  • We perform ablation studies and model attribution analysis to assess RainPro-8 performance and the effects of key design choices and input features.
@@ 2 Related Work
Traditional methods for weather forecasting are mostly numerical weather prediction (NWP) models, which simulate atmospheric dynamics using mathematical equations, also as ensembles over multiple simulations (Toth and Kalnay, 1997). NWP models demand significant compute, especially for ensembles, which restricts their spatial and temporal resolution. Their spin-up time results in poor performance for short lead times (Ma et al., 2021). Extrapolation-based methods, e.g. PySTEPS (Pulkkinen et al., 2019), RainyMotion (Ayzel et al., 2019) are limited to reduced forecast lengths due to their assumptions on constant motion and intensity (van Nooten et al., 2023).
Deep learning formulates precipitation nowcasting as a spatio-temporal prediction (An et al., 2024). Approaches in this area include the use of convolutions in recurrent networks (Shi et al., 2015, 2017; Wang et al., 2023b; Ma et al., 2024), the U-Net architecture (Ayzel et al., 2020; Fernández and Mehrkanoon, 2021; Zhang et al., 2023; Trebing et al., 2021), transformers (Gao et al., 2024b; Wu et al., 2024), diffusion models (Leinonen et al., 2023; Gao et al., 2024a; Gong et al., 2024), and adversarial training (Ravuri et al., 2021). Additionally, some research focuses on improving the loss function to tackle issues such as data imbalance, blurriness, and sparsity in precipitation nowcasting models (Xu et al., 2024a; Ko et al., 2022; Yan et al., 2024). Despite extensive research in this area, most efforts have focused on short-term forecasts (1 to 3 hours) using only radar, where the reduced lead time eliminates the need for multiple data sources, large contexts, or accounting for the higher uncertainty associated with 8-hour predictions.
Limited work exists in precipitation forecasting for lead times beyond 3 hours. The Weather4Cast competition (Gruca et al., 2022) tackles 8-hour precipitation forecasting from satellite alone, but it has shown limited success compared to traditional methods (Li et al., 2023) due to challenges in estimating current precipitation from satellite imagery and overcoming resolution issues compared to radar. For 6-hour precipitation nowcasting, NPM (Park et al., 2024) relies solely on satellite data without data fusion, while Nowcast-to-Forecast (An, 2023) combines satellite and radar but is limited by a biased rainy-day dataset. Kim et al. (2024) only compare hourly forecasts to extrapolation baselines, which are ineffective for longer-term predictions. TAFFNet (Wang et al., 2023a) provides 12-hour predictions but neglects uncertainty and relies on radar and NWP data at equal resolutions, which are not always available.
@@
Another line of work focuses on medium-range forecasting. Medium-range models like GraphCast (Lam et al., 2023), Pangu-Weather (Bi et al., 2023), or GenCast (Price et al., 2024) rely on reanalysis datasets like ERA5 (Hersbach et al., 2020), which suffer from low resolution, delays, and biases, especially in surface variables and precipitation (Lavers et al., 2022). Precipitation, highly variable and challenging to predict, is underrepresented in such models (Rasp et al., 2024), with most failing to integrate critical observational data like radar. While some models, like NeuralGCM (Yuval et al., 2024), incorporate global precipitation predictions, they lack the high resolution and operational frequency needed for our setting. Other work (Xu et al., 2024b) has tried to address the gap between nowcasting and medium-range forecasting, but model performance drops at high temporal resolutions and focuses only on low precipitation.
3 Method
@@ 3.1 Probabilistic Precipitation Forecasting
Radar data alone is not enough for accurate precipitation forecasting over long periods (e.g., up to 8 hours) because its reliance on ground-based systems limits its coverage and it cannot capture atmospheric conditions beyond water vapor. To enhance forecasting accuracy for longer lead times, additional data sources like satellite or NWP provide broader coverage and a more comprehensive representation of the atmosphere. Such additional sources introduce challenges due to differences in temporal frequency and spatial resolution, requiring careful preprocessing and alignment as detailed in Appendix A. The generalized input $X$, with heterogenous sources Sources, is given by:
@@
In addition to spatiotemporal accuracy, another critical priority is generating probabilistic forecasts that address the question: What is the probability of a specific amount of rainfall at a given location and time? Commonly, quantifying uncertainty involves generating multiple forecasts for ensembles to calculate probabilities, but this approach demands substantial computational resources. Instead of relying on ensemble methods, our approach directly predicts the probability distribution of precipitation intensities, similar to MetNet 1-3 (Sønderby et al., 2020; Espeholt et al., 2022; Andrychowicz et al., 2023). Our goal is to generate accurate probability maps to better capture uncertainty and variability in precipitation patterns, instead of radar-like outputs, which become less reliable over longer lead times. To model probabilities for different precipitation intensity classes $I$, we redefine the target as probability maps for each precipitation intensity class:
@@
In the case of MetNet, $P_{t,c}=P(R_{t}\in I_{c})$, which predicts the probability of precipitation intensity falling within specific predefined ranges. Given this definition, MetNet is trained with the cross entropy loss. However, this approach ignores the intrinsic order of the intensity classes, which we aim to solve with the proposed Ordinal Consistent loss.
@@ 3.2 Ordinal Consistent Loss
By having the model output $P(R_{t}\geq\min(I_{c})|R_{t}\geq\min(I_{c-1}))$ within the range $[0,1]$ for each intensity class $c$, we guarantee monotonic probabilities $P_{t,c}$ that respect the inherent order of precipitation classes. For the lowest intensity class, the model outputs $P(R_{t}\geq\min(I_{1})$. This approach, as proposed by Fernandes and Cardoso (2018) for semantic segmentation, improves both interpretability and consistency in the generated forecasts.
For any given class, the predicted probability $P_{t,c}$ is computed as the cumulative product of the model’s outputs for all preceding classes:
@@
The loss uses target binary masks indicating whether $R_{t}\geq\min(I_{c})$ for each class and timestep. Binary Cross Entropy (BCE) is used to compute the loss between the target masks and predicted probabilities. This calculation is performed without reduction to yield a per-pixel loss for each spatial location, class, and timestep.
To enhance training, the ordinal consistent mask (OC mask), computed for every sample as $R_{t}\geq\min(I_{c-1}))$, is incorporated. It ensures that for each intensity class $c$, the loss is averaged only over pixels where the previous class is activated, helping the model take advantage of the implicit ordinality. The loss is also not averaged over prediction pixels that lack ground truth values, typically missing due to the limited coverage of radar data. The loss function, based on BCE with the OC mask, is defined as:
$\mathcal{L}$ $=\text{mean}\left(\text{loss}(t,c,h,w)\right)\quad\text{if}\quad\text{mask}(t,c,h,w)\neq 1,$ (6)
[…]
$\quad\quad\quad\quad\text{mask}(t,c,h,w)=\mathds{1}\left(R_{t}(h,w)=-1\right)\lor\mathds{1}\left(R_{t}(h,w)<\min(I_{c-1})\right),$
where $R_{t}(h,w)=-1$ in areas without radar coverage.
@@ 3.3 Single-Pass Predictions
Instead, we make single-pass predictions to generate all forecast timesteps in one forward pass, reducing resource requirements, accelerating training convergence and inference, and improving temporal consistency in the outputs. The model encodes timestamps into the channel dimension, with the output $B(TC)HW$ reshaped to $B\,T\,C\,H\,W$, where $T$ represents all timesteps.
Inspired by MetNet-3’s lead-time sampling strategy, we introduce lead time weights. With lead-time sampling, lead times are sampled using a biased scheme based on an exponential distribution, which results in more frequent use of samples with shorter lead times during training, giving them a larger influence on model loss. In contrast, our approach includes all lead times in every training sample, so we incorporate lead time-dependent weights into the loss function to reduce the weighting of longer lead times. This ensures the model still prioritizes earlier lead times during training, improving performance across the entire range of lead times, including the more challenging longer ones. Therefore, the pixel-wise loss function is multiplied by lead-time weight $W_{t}$, normalized from an exponential distribution with a given decay rate $\alpha$, determining the relative weight assigned to the first timestep compared to the last one.
$\textrm{weights}_{\textrm{exp}}[t]=\exp(-\alpha\times t)\quad\textrm{weights}_{\textrm{norm}}[t]=\frac{\textrm{weights}_{\textrm{exp}}[t]}{\sum\textrm{weights}_{\textrm{exp}}[t]}\quad W_{t}=\frac{\textrm{weights}_{\textrm{norm}}[t]}{\overline{\textrm{weights}_{\textrm{norm}}}}\$ (7)

### v2

@@ Abstract
We present a deep learning model for high-resolution probabilistic precipitation forecasting over an 8-hour horizon in Europe, overcoming the limitations of radar-only deep learning models with short forecast lead times. Our model efficiently integrates multiple data sources - including radar, satellite, and physics-based numerical weather prediction (NWP) - while capturing long-range interactions, resulting in accurate forecasts with robust uncertainty quantification through consistent probabilistic maps. Featuring a compact architecture, it enables more efficient training and faster inference than existing models. Extensive experiments demonstrate that our model surpasses current operational NWP systems, extrapolation-based methods, and deep-learning nowcasting models, setting a new standard for high-resolution precipitation forecasting in Europe, ensuring a balance between accuracy, interpretability, and computational efficiency. Code is available at URL.
@@ 1 Introduction
Recent advances in artificial intelligence have generated significant interest in deep learning for weather forecasting (Rasp et al., 2024; An et al., 2024). Deep learning excels at handling complex, large-scale, high-dimensional data, making it well-suited for capturing intricate, nonlinear patterns in spatio-temporal systems (Manzhu Yu & Li, 2024). Although deep learning has achieved remarkable success in both nowcasting (Gao et al., 2024b; Gong et al., 2024) and medium-range forecasting (Lam et al., 2023; Price et al., 2024), significant challenges remain. Deep learning models for nowcasting are often limited to very short lead times (up to two hours). In contrast, medium-range models, which predict broader atmospheric dynamics for up to 10 days, typically operate at coarser resolutions and are influenced by precipitation-specific biases in the training datasets (Lavers et al., 2022). As a result, they struggle to capture small-scale precipitation features, like local showers, often leading to the exclusion of precipitation forecasts in medium-range models (Lam et al., 2023).
@@
This work addresses the challenge of forecasting precipitation for up to 8 hours at high spatio-temporal resolutions, bridging the gap between nowcasting and medium-range forecasting. Forecasting over an 8-hour horizon is critical for timely predictions that help mitigate risks like flooding and optimize resource management in agriculture, energy, or transportation. This task is particularly difficult due to the stochastic and sparse nature of precipitation, especially over extended lead times. Designing a deep learning model for this task requires addressing skewed and intermittent precipitation distributions, incorporating multi-sensor data, and ensuring robust uncertainty quantification.
@@
Most existing deep learning-based nowcasting methods focus on generating radar-like precipitation forecasts (Ravuri et al., 2021; Gao et al., 2024b), but these become increasingly challenging and less accurate with longer lead times. Instead, probabilistic models emphasize uncertainty estimation and provide a more accurate and practical representation of rainfall at varying intensities. The MetNet family of models (Sønderby et al., 2020; Espeholt et al., 2022; Andrychowicz et al., 2023) represents state-of-the-art deep learning-based systems for probabilistic precipitation forecasting. These models produce high-resolution forecasts for 8-24 hours in the United States, outperforming operational NWP systems. They achieve efficient probabilistic forecasting using cross-entropy loss on precipitation bins, requiring only one forward pass rather than an ensemble. However, this approach neglects the ordinality between bins. In addition, they rely on lead time conditioning, where the model generates forecasts for only one lead time at a time. While this approach is fully parallelizable across multiple GPUs, it significantly increases inference computational demands. Moreover, training MetNet requires substantial computational resources and is limited to the United States.
@@
We propose RainPro-8, an efficient deep learning model to estimate rainfall probabilities over 8 hours by determining the probability of different levels of precipitation at a given location and time. RainPro-8 is an efficient model based on MetNet-3 (Andrychowicz et al., 2023) that uses less than 20% of MetNet-3’s training parameters to achieve 8-hour high-resolution precipitation forecasting across Europe — a region characterized by diverse climates, complex terrain, and highly variable precipitation dynamics (Ehmele et al., 2020). Its training objective explicitly accounts for the ordinality between precipitation bins. Lastly, it demonstrates that it is possible to generate all lead times simultaneously by downweighing later lead times during training, greatly enhancing inference efficiency and improving temporal consistency. RainPro-8 model shows substantial improvements over traditional precipitation forecasting methods, including NWP models and extrapolation-based techniques, as well as deep-learning nowcasting models, across various rain intensities for the next 8 hours. Evaluation on the widely used SEVIR (Veillette et al., 2020) benchmark also demonstrates strong and robust performance beyond 8-hour multi-source tasks. Our main contributions include:
@@
  • Efficient architecture and training strategy for multi-source data of varying temporal or spatial resolution, consistent prediction of precipitation classes for all lead times at once.
  • Extensive empirical evaluation demonstrating that RainPro-8 outperforms existing operational methods by 65% and achieves significant gains over deep-learning nowcasting models; ablation studies and model attribution analysis of key design choices and model input.
  • Demonstration of RainPro’s versatility for radar-only 2-hour predictions on the SEVIR benchmark, achieving state-of-the-art performance compared to both deterministic and generative nowcasting models.
@@ 2 Related Work
Traditional methods for weather forecasting are mostly numerical weather prediction (NWP) models, which simulate atmospheric dynamics using mathematical equations, also as ensembles over multiple simulations (Toth & Kalnay, 1997). NWP models demand significant compute, especially for ensembles, which restricts their spatial and temporal resolution. Their spin-up time results in poor performance for short lead times (Ma et al., 2021). Extrapolation-based methods, e.g. PySTEPS (Pulkkinen et al., 2019), RainyMotion (Ayzel et al., 2019) are limited to reduced forecast lengths due to their assumptions on constant motion and intensity (van Nooten et al., 2023).
Deep learning formulates precipitation nowcasting as a spatio-temporal prediction (An et al., 2024). Approaches in this area include the use of convolutions in recurrent networks (Shi et al., 2015; 2017b; Wang et al., 2023b; Ma et al., 2024), the U-Net architecture (Ayzel et al., 2020; Fernández & Mehrkanoon, 2021; Zhang et al., 2023; Trebing et al., 2021), transformers (Gao et al., 2024b; Wu et al., 2024), diffusion models (Leinonen et al., 2023; Gao et al., 2024a; Yu et al., 2024; Gong et al., 2024), and adversarial training (Ravuri et al., 2021). Some work improves loss functions to tackle data imbalance, blurriness, and sparsity in precipitation nowcasting models (Xu et al., 2024a; Ko et al., 2022; Yan et al., 2024). Still, most efforts focus on short-term forecasts (1 to 3 hours) using only radar, where reduced lead time eliminates the need for multiple data sources, large contexts, or accounting for higher uncertainty associated with 8-hour predictions.
Limited work exists for lead times beyond 3 hours. The Weather4Cast competition (Gruca et al., 2022) tackles 8-hour precipitation forecasting from satellite alone, but it has shown limited success compared to traditional methods (Li et al., 2023) due to challenges in estimating current precipitation from satellite imagery and resolution issues compared to radar. For 6-hour precipitation nowcasting, NPM (Park et al., 2024) relies solely on satellite data without data fusion, while Nowcast-to-Forecast (An, 2023) combines satellite and radar but is limited by a biased rainy-day dataset. Kim et al. (2024) only compare hourly forecasts to extrapolation baselines, which are ineffective for longer-term predictions. TAFFNet (Wang et al., 2023a) provides 12-hour predictions but neglects uncertainty and relies on radar and NWP data at equal resolutions not always available in practice.
@@
Medium-range forecasting models like GraphCast (Lam et al., 2023), Pangu-Weather (Bi et al., 2023), or GenCast (Price et al., 2024) rely on reanalysis datasets like ERA5 (Hersbach et al., 2020), which suffer from low resolution, delays, and biases, especially in surface variables and precipitation (Lavers et al., 2022). Precipitation, highly variable and challenging to predict, is underrepresented in such models (Rasp et al., 2024), with most failing to integrate critical observational data like radar. While some models, like NeuralGCM (Yuval et al., 2024), incorporate global precipitation predictions, they lack the high resolution and operational frequency needed for our setting. Other work (Xu et al., 2024b) has tried to address the gap between nowcasting and medium-range forecasting, but model performance drops at high temporal resolutions and focuses only on low precipitation.
3 RainPro-8 Method
@@ 3.1 Probabilistic Precipitation Forecasting
Radar data alone is insufficient for accurate precipitation forecasting over 8 hours because its ground-based systems limit its coverage and cannot capture atmospheric conditions beyond water vapor. Additional data sources like satellite or NWP offer broader coverage and a more comprehensive representation of the atmosphere. Such additional sources introduce challenges due to differences in temporal frequency and spatial resolution, requiring careful preprocessing and alignment as detailed in Appendix A. The generalized input $X$, with heterogenous sources Sources, is given by:
@@
In addition to spatiotemporal accuracy, a critical priority are probabilistic forecasts: What is the probability of a specific amount of rainfall at a given location and time? Commonly, quantifying uncertainty involves generating multiple forecasts for ensembles to calculate probabilities, suffering from substantial computational demands. Instead of ensemble methods, our approach directly predicts the probability distribution of precipitation intensities, similar to MetNet 1-3 (Sønderby et al., 2020; Espeholt et al., 2022; Andrychowicz et al., 2023). Our goal is to generate accurate probability maps to capture uncertainty and variability in precipitation patterns, instead of radar-like outputs, which become less reliable over longer lead times. To model probabilities for different precipitation intensity classes $I$, we redefine the target as probability maps for each precipitation intensity class:
@@
In MetNet, $P_{t,c}=P(R_{t}\in I_{c})$ predicts the probability of precipitation intensity within predefined ranges, trained under cross entropy loss. However, this approach ignores the intrinsic order of the intensity classes, which we address with our proposed Ordinal Consistent loss.
@@ 3.2 Ordinal Consistent Loss
Model output $P(R_{t}\geq\min(I_{c})|R_{t}\geq\min(I_{c-1}))$ for each intensity class $c$ guarantees monotonic probabilities $P_{t,c}$ that respect the inherent order of precipitation classes. For the lowest intensity class, the model outputs $P(R_{t}\geq\min(I_{1}))$. This approach, as in Fernandes & Cardoso (2018) for semantic segmentation, improves both interpretability and consistency in the generated forecasts. For any class, probability $P_{t,c}$ is the cumulative product of model outputs for all preceding classes:
@@
The loss uses target binary masks indicating whether $R_{t}\geq\min(I_{c})$ for each class and timestep. The loss between the target masks and predicted probabilities is Binary Cross Entropy (BCE) without reduction to yield a per-pixel loss for each spatial location, class, and timestep. The ordinal consistent mask for every sample $(R_{t}\geq\min(I_{c-1}))$ ensures that for each intensity class $c$, the loss is averaged only over pixels where the previous class is activated, encouraging the model to take advantage of class ordinality. Note that the loss is not averaged over prediction pixels that lack ground truth values ($R_{t}(h,w)=-1$), typically missing due to the limited coverage of radar data.
$\mathcal{L}=\frac{1}{|S|}\sum_{(t,c,h,w)\in S}\operatorname{BCE}(t,c,h,w),\quad\text{where }S=\{(t,c,h,w):R_{t}(h,w)\geq\min(I_{c-1})\}$ (6)
@@ 3.3 Single-Pass Predictions
Instead, we make single-pass predictions to generate all forecast timesteps in one forward pass, reducing resource requirements, accelerating training convergence and inference, and improving temporal consistency in the outputs. The model encodes timestamps into the channel dimension, with the output $B(TC)HW$ reshaped to BTCHWB\,T\,C\,H\,W, where $T$ represents all timesteps.
MetNet-3 uses lead time sampling for more frequent use of samples with shorter lead times during training. In contrast, we propose lead time weights, where all lead times are included in every training sample, but are weighted in the loss function to reduce the impact of longer lead times. This ensures that the model still prioritizes earlier lead times during training, and improves performance across the entire range of lead times, including the more challenging longer ones. Thus, we multiply the pixel-wise loss function by lead time weight $W_{t}$, normalized from an exponential distribution with decay rate $\alpha$, determining the relative weight of the first timestep over the last one.
$\textrm{weights}_{\textrm{exp}}[t]=\exp(-\alpha\times t)\quad\textrm{weights}_{\textrm{norm}}[t]=\frac{\textrm{weights}_{\textrm{exp}}[t]}{\sum\textrm{weights}_{\textrm{exp}}[t]}\quad W_{t}=\frac{\textrm{weights}_{\textrm{norm}}[t]}{\overline{\textrm{weights}_{\textrm{norm}}}}\\$ (7)